Operations Report – April 18th

Crew 192 Operations Report 18Apr2018

SOL 10

Name of person filing report: Victoria Varone

Non-nominal systems: None

Generator (hours run): 12 hours

Solar ‒ SOC 66%

Diesel – 75% fuel

Propane – 62% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – All ATVs filled 100%, 25% left in one can, requesting gas refill.

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 690 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133669.6

Toilet tank emptied: No. Last emptied April 16.

ATV’s Used: None

Oil Added? No but checked and OK.

ATV Fuel Used: 0% today

# Hours the ATVs were Used today: 0

Notes on ATVs: ATVs fine, oil checked, gas filled yesterday.

Deimos rover used: Yes (out of commission per Shannon)
Hours: 121.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes
Hours: 32.8 (corrected)
Beginning charge: 100%
Ending charge: 68%
Currently charging: Yes

Opportunity rover used: yes
Hours: 29.8
Beginning charge: 100%
Ending charge: 81%
Currently charging: Yes

Curiosity rover used: Yes
Hours: 27.4
Beginning charge: 100%
Ending charge: 67%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Some great EVAs on a beautiful day.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, three one-piece suits used, radios working fine.

Summary of Hab operations: Cooked, cleaned, and organized the hab.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: None

Summary of RAM operations: Trailer water tank remains in the RAM to air out after being cleaned for the final time yesterday, will be checked tomorrow.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: – Hab upper deck faucet hot water shut due to increasing leaks.

– Fridge got hot last night despite being closed and plugged in properly, was unplugged and plugged back and is working fine again.

– Steps in front of rear airlock were secured again after become a bit loose after the first fix.

Copyright © The Mars Society. All rights reserved. | Main Site